if y = 3+ 3^1/3 +3^2/3 then the value of y^3 - 9y^2+18y

Grade:10

1 Answers

Probin pegu
6 years ago
Y =3+ 3^1/3 +3^2/3 = 7Now, y3+9y2 - 18y=7^3+9×7^2+ 18×7 = 343 +9×49 +126 = 343+ 441+126 = 910Therefore the value of 3+ 3^1/3 +3^2/3 is 910